40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14110     Entry time: Sat Jul 28 00:45:11 2018
Author: terra, sandrine 
Type: Summary 
Category: Thermal Compensation 
Subject: Heater measurements overview 

[Sandrine, Koji, Terra]

Summary: We completed multiple scans at different heating powers for the reflector set up, observing unique HOM peak shifts of tens of kHz. We also observed HOM5 shifts with the cylinder set up. Initial Lorentzian fittings of the magnitude give tens of Hz resolution. I summarize the main week's work below. 

Set-up

Heater set-up is described in several previous elogs, but attachments #1 and #2 show the full heater set-up and wiring/pinouts in and out of vacuum, since we're all intimately aware of how confusing in-vacuum pinouts can be. We are not using the Sorenson power supply (as described in 14071); we just have the BKPrecision power supply 1735 sitting next to the ETMY rack and are manually going out to turn on/off. 

We've continued to use the scan setup described in elog 14086, which is run using /users/annalisa/postVent/AGfast.py. Step by step notes for setting up the scan, running the scans, and processing the scans are attached in notes.txt.

Inducing/witnessing HOMs

The aux input beam was already clipped and on wednesday (after Trans was centered, 14093) we also clipped the output aux beam with razor blade (angled vertically and horizontally, elog 14103) before PDA255; we clipped ~1/3 of the output beam. Attachment #3 shows before and after clipping output, where orange 'cold' == unclipped, black 'mean' == clipped (all in cold state). Up to HOM5 is visible. 

Measurements

Below is a summary of the available scan data. We also have cold (0A) scans CAR-HOM5 and full FSR scans for most configurations. 

Elliptic Reflector
current[A] voltage[V] power[W] scans
0.4 2 0.8 CAR-HOM3(x1)
0.5 3.4 1.7 CAR-HOM3(x1)
0.6 5 3.0 CAR-HOM3(x1)
0.8 9.4(9.7) 7.5(7.8) CAR-HOM5(>x5)
0.9 12 10.8 CAR-HOM5(x4)
1.09 17 18.5 CAR-HOM3

 

 

 

 

 

 

 

Cylinder + Lenses
current[A] voltage[V] power[W] scans
0.9 15 13.5 CAR-HOM5(odds x4)

We tried the cylinder set-up again tonight for the first time since inital try and can see shifts of HOM5 - see attachment #5; we haven't looked in detail yet, but it looks like odd modes are more effected, suggesting the ring heat pattern is off centered from the beam axis. 

Scan data is saved in the following format: users/annalisa/postVent/scandata/{reflector,cylinder}/{parsed,unparsed}/{CAR,HOM1,HOM2,HOM3,HOM4,HOM5}{_datetime}{_parsed,_unparsed}.{txt,pdf}

Minimum heating

On 7/26 we increased the power to the elliptical reflector heater in steps to find the minimum heater power required to see frequency shifts with our measurement setup. Lowest we can resolve is a shift in HOM3 with 1.7W (0.5A/3.4V). According to Annalisa's measurements in elog 14050, this would be something like 30-60 mW radiated power hitting the test mass. We only looked at CAR - HOM3 for this investigation; data for scans at 0.4A, 0.5A, 0.6A is available as indicated above.

Lorentizian Fitting

The Lorentzian fitting was done using the equation a + b / sqrt(1+((x-c)/d*2), where a = constant background, b = peak height above background, c = peak frequency, d = full width at half max. 

The fitting is still being edited and optimized. We will crop the data to zoom in around the peak more.

The Lorentzian fit of the magnitude shows ~10Hz of resolution. (See attachment 6 for the carrier at 8A and attachment 7 for HOM 1 at 9A)

We're working on fitting the full complex data.

 

 

Attachment 1: heater_setup.jpg  26 kB  Uploaded Sat Jul 28 02:12:54 2018  | Hide | Hide all | Show all
heater_setup.jpg
Attachment 2: heater_wiring.jpg  19 kB  Uploaded Sat Jul 28 02:41:08 2018  | Hide | Hide all | Show all
heater_wiring.jpg
Attachment 3: notes.txt  1 kB  Uploaded Sat Jul 28 03:13:54 2018  | Hide | Hide all | Show all
Notes for running scans:
1. when first turning on Agilent, set initial stuff
    > cd /users/annalisa/postVent/20180718
    > AGmeasure TFAG4395Atemplate.yml
2. tweak arm alignment and offset PLL
    > sitemap (then IFO --> ALIGN and also PSL --> AUX)
    > to increase 
3. make sure X-arm is misagligned (hit '! Misalign' button for ITMX, ETMX) 
3. run scan
    > python AGfast.py startfreq stopfreq points
... 36 more lines ...
Attachment 4: FSR_clipped.pdf  122 kB  Uploaded Sat Jul 28 03:14:19 2018  | Show | Hide all | Show all
Attachment 5: cylinderHOM5.pdf  74 kB  Uploaded Sat Jul 28 04:44:25 2018  | Hide | Hide all | Show all
cylinderHOM5.pdf
Attachment 6: pt8A_CAR.pdf  75 kB  Uploaded Sat Jul 28 05:02:26 2018  | Hide | Hide all | Show all
pt8A_CAR.pdf
Attachment 7: pt9A_HOM1.pdf  59 kB  Uploaded Sat Jul 28 05:02:44 2018  | Show | Hide all | Show all
ELOG V3.1.3-